OpenStudy (hellokitty17):

Solve 3x2 + 18x + 15 = 0 by completing the square. 3(x + 3)2 = 6; x = −3 + square root 2, x = −3 − square root 2 3(x + 9)2 = 12; x = −11, x = −7 3(x + 3)2 = 12; x = −5, x = −1 3(x + 9)2 = 228; x = −9 + square root 76, x = −9 − square root 76

OpenStudy (anonymous):

3x^2 + 18x + 15 = 0 3(x^2 + 6x + 5) = 0 3(x + 3)^2 - 12 = 0 Answer choice: c.) 3(x + 3)2 = 12; x = -5, x = -1

OpenStudy (hellokitty17):

thanks!

oregonduck (oregonduck):

yup

oregonduck (oregonduck):

source for @jcoury https://answers.yahoo.com/question/index?qid=20140327120313AA8hJ7M

OpenStudy (anonymous):

yup

oregonduck (oregonduck):

lol :)

OpenStudy (anonymous):

like i said: if i don't have to, i won't If i have to, i will